# Don’t Retrieve, Navigate: Distilling Enterprise Knowledge into Navigable Agent Skills for QA and RAG

### 16 Apr 2026 v1

@@ 5 Limitations
The per-query cost of Corpus2Skill is dominated by input tokens: the Skills API includes navigation file content in every call, so even relatively small SKILL.md and INDEX.md files accumulate significant token counts across turns. Corpus2Skill is best suited for high-value queries where answer quality justifies the cost premium, such as complex multi-hop support questions or compliance-critical lookups. For high-volume, low-stakes queries, single-shot retrieval or RAPTOR remain cost-effective. Prompt caching could substantially reduce this overhead by reusing repeated skill file content across queries within the same session.
@@ 7 Conclusion
Corpus2Skill demonstrates that giving an LLM agent a structured, browsable overview of a document corpus leads to higher-quality answers than providing it with retrieved passages alone. The key insight is that modern LLM agents are capable navigators when given the right map: a pre-compiled hierarchy of summaries and topic clusters enables informed, top-down exploration that exploits the agent’s reasoning abilities rather than treating it as a passive consumer of search results. Because the agent reads summaries describing each branch, it can make informed routing decisions and combine evidence across branches when queries span multiple topics—an advantage that flat retrieval, which offers no corpus overview, cannot match. While the approach incurs higher per-query cost due to navigation overhead, it achieves the highest quality on WixQA and suggests a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than at query time to make it searchable. Hierarchy depth grows logarithmically ($O(\log_{p}N)$), so the framework scales to larger corpora.
Future work includes evaluation on more domains, incremental compilation, and prompt caching to cut query cost.

### 29 Apr 2026 v2

@@ 5 Limitations
The per-query cost of Corpus2Skill is dominated by input tokens: the Skills API includes navigation file content in every call, so even relatively small SKILL.md and INDEX.md files accumulate significant token counts across turns. Enabling Anthropic ephemeral prompt caching on the navigation context reduces \$/query from \$0.172 to \$0.089—close to the Agentic baseline—since roughly 70% of the prompt content is reused across consecutive turns and queries; we use this cached configuration throughout our main results. Corpus2Skill is nonetheless best suited for high-value queries where answer quality justifies the cost premium, such as complex multi-hop support questions or compliance-critical lookups. For high-volume, low-stakes queries, single-shot retrieval or RAPTOR remain cost-effective.
@@
Corpora where corpus navigation does not help.
Our RAGBench generalization study (Section 4.4) makes the failure modes explicit. Corpus2Skill loses to flat retrieval on three structurally distinct corpus shapes. (1) Long extractive documents: CUAD legal contracts ($\approx$7,500 characters per doc, with answers being verbatim quotes from a single clause) and TechQA technotes (3,740 characters covering multiple symptom categories) violate the assumption that one document fits cleanly under one branch. (2) Homogeneous tabular corpora: TatQA is essentially identical SEC-style financial-statement tables, so the level-1 cluster summaries collapse onto near-duplicate labels and the navigator cannot disambiguate top-level branches; this case yields the worst quality and cost in our evaluation. (3) Open-domain pools where retrieval recall is the bottleneck: on CovidQA short research abstracts and ExpertQA cross-disciplinary scientific QA, hierarchical navigation adds turns and tokens without improving over what dense retrieval already surfaces in a single shot. We see these as honest scope limits of the design hypothesis behind Corpus2Skill rather than implementation issues. Three natural follow-on directions are clear from this analysis: a span-level retrieval head attached at the leaf level so the agent can quote verbatim from long documents (addressing CUAD), table-aware schema clustering that groups financial-statement rows by line-item / period / segment rather than embedding similarity (addressing TatQA), and hybrid routing that lets a lightweight classifier decide per-query whether to enter the skill tree or fall back to dense retrieval. We are exploring this last design under the working name CorpusForge in subsequent work.
@@ 7 Conclusion
Corpus2Skill demonstrates that giving an LLM agent a structured, browsable overview of a document corpus leads to higher-quality answers than providing it with retrieved passages alone. The key insight is that modern LLM agents are capable navigators when given the right map: a pre-compiled hierarchy of summaries and topic clusters enables informed, top-down exploration that exploits the agent’s reasoning abilities rather than treating it as a passive consumer of search results. Because the agent reads summaries describing each branch, it can make informed routing decisions and combine evidence across branches when queries span multiple topics—an advantage that flat retrieval, which offers no corpus overview, cannot match. With Anthropic ephemeral prompt caching enabled on the navigation context, Corpus2Skill matches the Agentic baseline on per-query cost (\$0.089 vs. \$0.088) while still delivering the highest quality on every metric, and it suggests a complementary design point for knowledge-grounded systems: invest computation at compile time to make the corpus navigable, rather than at query time to make it searchable. Hierarchy depth grows logarithmically ($O(\log_{p}N)$), so the framework scales to larger corpora.
Future work includes evaluation on more domains, incremental compilation, and multi-parent soft clustering to handle queries that span multiple topics.